# SpecExec: Massively Parallel Speculative Decoding for Interactive LLM Inference on Consumer Devices

### v2

@@ 4.3 Implementation details
Finally, there are several important technical improvements that speed up the real-world inference speed. When running the forward pass with an offloaded target model, we accelerate inference by loading the next layer parameters in parallel with computing the previous layer activations using a dedicated CUDA stream, which is known to speed up offloading in other use cases [16, 17, 19]. We also preload the first few layers of the target model on the GPU in the background while drafting for a further speedup. We describe additional implementation details in Appendix C.
In our experiments, we also consider quantizing target models using recent post-training quantization algorithms [28, 30, 29]. While quantization is generally popular among LLM practitioners, it is particularly useful for our use case, as quantized models take less time to load from RAM to GPU and have RAM offloading requirements attainable by consumer computers.
@@ 5.1 Probability Coverage
The core assumption behind Algorithm 1 is that a reasonably large draft can “cover” most of the high probability sequences of the target model. This is only possible if the target model predictions have low entropy (i.e., there is a small number of tokens with high probability) and the draft model can guess these tokens most of the time.
To test these assumptions in isolation, we measure the total probability mass “covered” by $k$ most likely tokens, as well as the probability of top-$k$ tokens guessed by draft models of varying size. If a certain draft model achieves a coverage probability $p$ for $k$ tokens, this entails taking the $k$ most likely tokens predicted by the draft model and measuring their probabilities with the main model (Llama-2-Chat 70B) would result in average cumulative probability equal $p$. We evaluated multiple draft models of various size: JackFram-160M [5], TinyLlama-1.1B-Chat v1.0 [33], Llama-2-Chat 7B and 13B [21]. We report these coverage probabilities on a sample of 512 OpenAssistant conversations [34]. For each conversation, we generate 64 additional tokens by sampling from the target 70B model probabilities. We sample these tokens using raw probabilities (without temperature or top-p sampling) and use the same tokens for every draft model.
The resulting coverage is reported in Figure 2. This leads to several important observations. First, the target model (Llama-2-Chat 70B) tends to have sharp probability distributions, where the top 1–4 tokens cover 90–98% of the entire probability mass. This agrees with existing observations that language models (especially the larger ones) are overconfident [35, 36].
Next, we compare how effective the draft models are at predicting these high-probability tokens. While all models eventually get over 90% coverage rate, Llama-2-Chat 7B makes much more accurate predictions with the first 1–4 tokens. This is important for our use case because, while the full draft tree contains thousands of tokens, individual tokens within that tree have much fewer children. Curiously, the 13B draft model demonstrates roughly the same accuracy as 7B despite its larger size.
Though we evaluate coverage for “raw” probabilities from the 70B model, many practical inference scenarios use temperature or nucleus sampling. In fact, the official default sampling method for Llama-2 70B uses both a temperature of 0.6 and top-0.9 nucleus sampling [37]. Generating in this way makes the model even more confident, which further improves the efficiency of parallel decoding.
@@ 5.2 Draft Acceptance Rates
Next, we study how speculative execution compares to existing speculative decoding variants for different token budgets. Since all algorithms guarantee that the tokens are sampled from $P(x_{t+1}|x_{0:t},\theta_{main})$, we compare them in ability to generate longest sequences of accepted given the same budget of draft tokens.
Since we are interested in very large budgets, we choose baseline algorithms that better fit this task. The original speculative decoding algorithm [3] generates a single sequence, which is truncated as soon as the algorithm rejects a single token. In other words, using a single long draft sequence results in most of the draft budget being wasted. Therefore we choose as baseline the SpecInfer algorithm that shares the large token budget across multiple stems in a tree.
Figure 3: Generation rate vs draft size for Llama 2-7B/70B chat models, MTBench [38] dataset.
Similarly to the previous section, we use the 70B Llama-2 and Llama-2-Chat as target models. The draft model choice was driven by both speed and acceptance rate factors: we found that using 7B draft models results in significantly more accepted tokens, and longer forward pass time is still affordable in the offloading setting. We report the effects of these draft models in more detail in Appendix D.
In each setup, we compared SpecExec and SpecInfer, using the 7B draft model, chosen based on Section 5.1. Figure 3 reports the average number of accepted tokens both for the default sampling configuration (temperature 0.6, top-p 0.9) and for greedy sampling for Llama2-70b-chat model using MTBench dataset. Similar tests were run for non-chat models on the C4 dataset, see Figure 4.
For smaller draft budgets, SpecExec performs on par with SpecInfer, but eventually outscales it as we increase the number of draft tokens. We attribute this not to the general verification algorithm, but to the fact that SpecExec constructs its draft out of most likely tokens, while SpecInfer must sample draft tokens to guarantee correctness. We also observe that speculative execution achieves a higher margin of improvement on MTBench samples than on C4. It also accepts more tokens with lower temperature. We attribute this to sharper token probability distributions, and thus, higher coverage rates for the same number of draft tokens.
Figure 4: Generation rate vs draft size for Llama 2-7B/70B models, C4 dataset.
@@ 5.3 Inference Speed
Finally, we evaluate the practical inference speed of SpecExec by running it with offloading in different hardware configurations. We run these evaluations for Llama 2 70B [21] models, both in regular and chat (instruction-tuned) versions. For prompts, we used subsamples of size 100 from OpenAssistant conversations [34], WikiText-2 [39], MTBench [38] and C4 [40], measuring the speed of generating 32+ tokens per prompt. For Llama 2, we tested two setups: running the main model in 16 bits or quantizing it to 4 bits using GPTQ [28]. We also tested Mixtral 8x7B [22] and Llama-3 [1] target models in fewer setups.
We measure the inference speed on multiple GPU types: A100 (high-end data-center), RTX 4090 (current generation high-end consumer), RTX 3090 (previous generation consumer), and RTX 2080Ti (older consumer GPU). The first three GPUs are connected to the host via PCIe Gen 4 x16, while 3090 and 2080Ti were tested with PCIe Gen 3 x16. Note that for A100, we report the forward pass time with offloading, even though the GPU can fit a quantized model in its memory. We run all experiments with a batch size of 1 to match the setup of running LLMs on a local machine.
Table 1: Inference speed with RAM offloading, A100 GPU, Chat / Instruct models, using SpecExec(SX) and SpecInfer(SI) methods
@@
Mistral-7b / Mixtral-8x7b-GPTQ OAsst 0.6 SX 256 12.38 3.58 3.5x
Llama3-8b / 70b OAsst 0.6 SX 1024 18.88 2.62 15.6x
@@
The average inference speed (measured in tokens per second) is reported in Tables 1 and  2. While the exact inference speed differs from setup to setup, Speculative Execution consistently speeds up generation with offloading by several times. These results compare favorably with recently published speculative decoding methods using fixed trees like Sequoia [41], which attains 2.2 tokens per second in the Llama 3 8B/70B setup, compared to 2.8 tokens per second in case of SpecExec.
In Table 3, we report similar experiments for a range of real-world consumer GPUs. Note that while the fastest inference time is on RTX 4090, slower consumer GPUs (e.g. RTX 2080Ti) still generate tokens quickly enough for interactive use. To fit into consumer setup we replaced a 16-bit Llama-2 70B model with 4-bit GPTQ compressed variant of Llama2-70B as target. To lower the VRAM requirements for 2080 Ti we used Sheared-LLaMA-1.3B [42] as a draft model, making the whole experiment consume just over 7Gb of VRAM.
Table 2: Inference speed with RAM offloading. A100 GPU, base models, using SpecExec(SX) and SpecInfer(SI)
@@
Mistral-7b / Mixtral-8x7b-GPTQ WikiText-2 0.6 SX 128 6.56 3.23 3.2x
Table 3: SpecExec inference on consumer GPUs with offloading, chat/instruct models, Llama-2-70B-GPTQ target model, $t=0.6$, OpenAssistant dataset.
@@
While this was not the primary focus area of our study, the SpecExec method can also deliver competitive speedups in inference without offloading. See Appendix F for sample results.
@@ 6 Conclusion and Future Work
In this work, we propose a method for fast inference of large models on consumer GPUs that unites the efficiency of offloading and speculative decoding in the large-token setup. The resulting method, SpecExec, shows competitive performance in real-world experimental setups, demonstrating the possibility of running large models locally at the speed of interactive inference.
@@ Appendix A Equivalence of Optimal Tree Search to Shortest Path Search
Here, $x_{i}\in\tau$ refers to a text token within the draft tree, $\mathcal{T}^{K}$ is a set of all trees of $K$ tokens and $P_{\text{reach}}(x_{i}|\tau)$ is the probability that the Speculative Execution verification phase accepts the full prefix $x_{0},\dots,x_{i-1}$ along the draft tree and considers sampling $x_{i}$ next. Finally, $\cdot P_{\text{accept}}(x_{i}|\tau)$ is the probability that the token $x_{i}$ will be accepted if it is reached during verification. Both $P_{\text{reach}}$ and $p_{\text{accept}}$ depend on the target model probabilities $P(x_{t+1}|x_{0:t},\theta_{\text{target}})$, which cannot be accessed in the drafting phase. Instead, we use the draft model to approximate the target model probabilities as follows:
@@
Since token probabilities cannot be greater than 1, the cumulative probability of $\pi(x_{i},\tau)\oplus x_{i}$ cannot exceed the cumulative probability of all tokens in $\pi(x_{i},\tau)$. Therefore, if a token $x_{i}$ is among the $K$ most likely tokens, every token in $\pi(x_{i},\tau)$ is also a part of the solution. Using this property, we can simplify the problem (3) as finding top-$K$ most likely prefixes, since they are guaranteed to form a tree. Formally speaking, the optimal tree consists of K tokens with the highest cumulative probability:
@@
This is similar (but not equivalent) to the standard beam search algorithm for neural sequence models [43, 44, 45]. The main difference is that beam search looks for complete sequences, while we need a tree of partial drafts. However, using beam search instead of solving Equation (3) directly leads to suboptimal drafts (see Appendix E for details).
Instead, we solve Equation (4) by reformulating it as a special case of shortest path problem. More specifically,
@@
In the worst case, this algorithm still makes up to $K$ steps if the solution to Equation (3) is a single “stem” B tokens long. However, the actual number of steps is significantly lower, often slightly above the lower bound $\lceil B/K\rceil$. In the practical GPU implementation, we also limit the maximum depth with a parameter $D$. The purpose of D is to limit the edge case where the draft model is very confident about the next token, and thus the solution to (3) is a single sequence of length K. For this edge case, Algorithm 2 will take long to generate a sequential draft, most of which will later be discarded if the draft model makes even one mistake.
Appendix B Parallel Graph Search
@@
There are dozens of works that study practical implementations of parallel shortest path search and SSSP in particular. One line of work proposes inexact search algorithms that use an approximate priority queue to improve the performance of SSSP: [46] proposes a queue with integer metric, [47] adds bucket fusion to reduce synchronization overhead..
@@
A significant effort was dedicated to efficient shortest-path search on GPUs, among others. [51] proposes a GPU-efficient SSSP that outperforms sequential CPU computations. [52, 53] compare several SSSP variants for GPUs. [54] adapts priority queues to run efficiently on GPU and uses the resulting data structure to accelerate SSSP.
Many works on parallel graph search focus on the distributed setting [55, 56, 57], addressing communication and synchronization overheads.
Finally, a large body of work studies the theoretical properties of parallel SSSP, including [58, 59, 60, 61, 62, 63, 64, 65].
Appendix C Additional Implementation Details

### v3

@@ 4.3 Implementation Details
Finally, we leverage several important technical improvements that speed up inference in real-world conditions. When running the forward pass with an offloaded target model, we accelerate inference by loading the next layer parameters in parallel with computing the previous layer activations using a dedicated CUDA stream, which is known to speed up offloading in other use cases Pudipeddi et al. (2020); Ren et al. (2021); Aminabadi et al. (2022). We also preload the first few layers of the target model on the GPU in the background while drafting for a further speedup. We describe additional implementation details in Appendix D.
In our experiments, we also consider quantizing target models using recent post-training quantization algorithms Frantar et al. (2022); Lin et al. (2023); Dettmers et al. (2023). While quantization is generally popular among LLM practitioners, it is particularly useful for our use case, as quantized models take less time to load from RAM to GPU and have RAM offloading requirements attainable by consumer hardware.
@@ 5.1 Probability Coverage
The core assumption behind Algorithm 1 is that a reasonably large draft can “cover” most of the high-probability sequences of the target model. This is only possible if the target model predictions have low entropy (i.e., there is a small number of tokens with high probability) and the draft model can guess these tokens most of the time.
To test these assumptions in isolation, we measure the total probability mass “covered” by $k$ most likely tokens, as well as the probability of top-$k$ tokens guessed by draft models of varying size. If a certain draft model achieves a coverage probability $p$ for $k$ tokens, this means that taking the $k$ most likely tokens predicted by the draft model and measuring their probabilities with the main model (Llama-2-Chat 70B) would result in an average cumulative probability equal to $p$. We evaluated multiple draft models of various size: JackFram-160M Miao et al. (2023), TinyLlama-1.1B-Chat v1.0 Zhang et al. (2024b), Llama-2-Chat 7B and 13B Touvron et al. (2023). We report these coverage probabilities on a sample of 512 OpenAssistant conversations Köpf et al. (2023). For each conversation, we generate 64 additional tokens by sampling from the target 70B model probabilities. We sample these tokens using original probabilities (without temperature or top-p sampling) and use the same tokens for every draft model.
The resulting coverage is reported in Figure 2. This leads to several important observations. First, the target model (Llama-2 Chat 70B) tends to have sharp probability distributions, where the top 1–4 tokens cover 90–98% of the entire probability mass. This agrees with existing observations that language models (esp. the larger ones) are overconfident Miao et al. (2021); Chen et al. (2023b).
Next, we compare how effective the draft models are at predicting these high-probability tokens. While all models eventually get over 90% coverage rate, Llama-2 Chat 7B makes much more accurate predictions with the first 1–4 tokens. This is important for our use case because, while the full draft tree contains thousands of tokens, individual tokens within that tree have much fewer children. Curiously, the 13B draft model demonstrates roughly the same accuracy as 7B despite its larger size.
Though we evaluate coverage for “raw” probabilities from the 70B model, many practical inference scenarios use temperature or nucleus sampling Holtzman et al. (2020). In fact, thedefault generation parameters for Llama-2 70B use both a temperature of 0.6 and top-0.9 nucleus sampling Face (2024). Generating in this way makes the model even more confident, which further improves the efficiency of parallel decoding.
@@ 5.2 Draft Acceptance Rates
Next, we study how Speculative Execution compares to existing speculative decoding variants for different token budgets. Since all algorithms guarantee that the tokens are sampled from $P(x_{t+1}|x_{0:t},\theta_{main})$, we compare them in their ability to generate longest sequences of accepted tokens given the same budget of draft tokens.
Since we are interested in very large budgets, we choose baseline algorithms that better fit this task. The original speculative decoding algorithm Leviathan et al. (2023) generates a single sequence, which is truncated as soon as the algorithm rejects a single token. In other words, using a single long draft sequence results in most of the draft budget being wasted. Therefore, as our baseline, we choose the SpecInfer algorithm that shares the large token budget across multiple stems in a tree.
Figure 3: Generation rate depending on the draft budget size for Llama 2-7B Chat as the draft model and Llama 2-70B Chat as the target model, MTBench Zheng et al. (2023) dataset. Results are obtained with an A100 GPU.
Similarly to the previous section, we use 70B versions of Llama 2 and Llama 2-Chat as target models. The draft model choice was driven both by the speed and the acceptance rate factors: we found that using draft models with 7B parameters results in significantly more accepted tokens, and a longer forward pass time is still affordable in the offloading setting. We report the effects of these draft models in more detail in Appendix E.
In each setup, we compared SpecExec and SpecInfer, using the 7B draft model, chosen based on our findings from Section 5.1. Figure 3 reports the average number of accepted tokens both for the default sampling configuration (temperature 0.6, top-p 0.9) and for greedy sampling for Llama 2-70B Chat model using the MTBench dataset. Similar tests were run for non-chat models on the C4 dataset; see Figure 4 for results.
Figure 4: Generation rate depending on the draft budget size for Llama 2-7B Chat as the draft model and Llama 2-70B Chat as the target model, C4 dataset. Results are obtained with an A100 GPU.
For smaller draft budgets, SpecExec performs on par with SpecInfer, but eventually outscales it as we increase the number of draft tokens. We attribute this not to the general verification algorithm, but to the fact that SpecExec constructs its draft out of most likely tokens, while SpecInfer must sample draft tokens to guarantee correctness. We also observe that Speculative Execution achieves a higher margin of improvement on MTBench samples than on C4. It also accepts more tokens with a lower temperature. We attribute this to sharper token probability distributions, leading to higher coverage rates for the same number of draft tokens.
@@ 5.3 Inference Speed
Finally, we evaluate the practical inference speed of SpecExec by running it with offloading in different hardware configurations. We run these evaluations for Llama 2-70B Touvron et al. (2023) models, both in regular and chat (instruction-tuned) versions. For prompts, we used subsamples of size 100 from OpenAssistant conversations Köpf et al. (2023), WikiText-2 Merity et al. (2016), MTBench Zheng et al. (2023), and C4 Raffel et al. (2020), measuring the speed of generating 32+ tokens per prompt. For Llama 2, we tested two setups: running the main model in 16 bits or quantizing it to 4 bits using GPTQ Frantar et al. (2022). We also tested Mixtral 8x7B Jiang et al. (2024) (also quantized with GPTQ) and Llama 3 AI (2024) target models in fewer setups.
Table 1: Inference speed with RAM offloading, A100 GPU, Chat / Instruct models, using SpecExec (SX) and SpecInfer (SI) methods. Generation rate (“Gen. rate”) denotes the average number of draft model tokens accepted for one target model iteration.
@@
Mistral-7B / Mixtral-8x7B OAsst 0.6 SX 256 12.38 3.58 3.5x
Llama 3-8B / 70B 0.6 SX 1024 18.88 2.62 15.6x
@@
We measure the inference speed with multiple GPU types: A100 (data-center GPU), RTX 4090 (current generation high-end consumer GPU), RTX 3090 (previous generation consumer GPU), and RTX 2080Ti (older consumer GPU). The first three GPUs are connected to the host via PCIe Gen 4 x16, while 3090 and 2080Ti were tested with PCIe Gen 3 x16. Note that for A100, we report the forward pass time with offloading, even though the GPU can fit a quantized model in its memory. We run all experiments with a batch size of 1 to match the setup of running LLMs on a local machine.
The average inference speed (measured in tokens per second) for A100 GPUs is reported in Tables 1 and  2. While the exact inference speed differs from setup to setup, Speculative Execution consistently speeds up generation with offloading by several times. These results compare favorably with recently published speculative decoding methods using fixed trees like Sequoia Chen et al. (2024b), which attains 2.2 tokens per second in the Llama 3-8B/70B setup, compared to 2.8 tokens per second in case of SpecExec.
In Table 3, we report the results of similar experiments for a range of real-world consumer GPUs. To reduce the memory requirements for the consumer setup, we replaced a 16-bit Llama-2 70B model with a 4-bit GPTQ compressed variant of Llama-2-70B as the target model. To lower the VRAM requirements for 2080 Ti, we used Sheared-Llama-1.3B Xia et al. (2024b) as a draft model, making the whole experiment consume just over 7 GB of VRAM. Note that while the fastest inference time is achieved on RTX 4090, slower consumer GPUs (for example, RTX 2080Ti) still generate tokens quickly enough for interactive use.
Table 2: Inference speed with RAM offloading. A100 GPU, base models, using SpecExec (SX) and SpecInfer (SI). Generation rate (“Gen. rate”) denotes the average number of draft model tokens accepted for one target model iteration.
@@
Mistral-7B / Mixtral-8x7B WikiText-2 0.6 SX 128 6.56 3.23 3.2x
Table 3: SpecExec inference speed on consumer GPUs with offloading, chat/instruct models, Llama 2 70B-GPTQ target model, $t=0.6$, OpenAssistant dataset.
@@
We also explore the relationship between the inference speed and the draft tree size. A larger draft budget allows for a greater number of tokens to be generated per step (see Figure 5 (left)). However, beyond a certain size threshold (hundreds or thousands of tokens, depending on the model and GPU), the time required for generation increases at an accelerating rate (see Figure 1 (right)). Consequently, the optimal draft tree size is typically smaller than the size that maximizes the token acceptance rate. According to our findings, displayed in Figure 5 (right), the optimal draft tree size is 128–512 for SpecInfer and 1024–2048 for SpecExec for the A100 GPU.
Figure 5: Acceptance counts (left) and generation speed (right) depending on the draft size. Llama 2-7B is used as the draft model, offloaded Llama 2-70B is the target model, MTBench dataset, t=0.6 and top-p=0.9. Results are obtained with an A100 GPU.
While this was not the primary focus area of our study, the SpecExec method can also deliver competitive speedups in inference without offloading; see Appendix G for sample results. Additional tests of SpecExec in generation with penalties show that the method is robust with such conditions: Appendix H provides the results of such an evaluation.
@@ 6 Conclusion and Future Work
In this work, we propose a method for fast inference of large models on consumer GPUs that unites the efficiency of offloading and speculative decoding in the large-budget setup. The resulting method, SpecExec, shows competitive performance in real-world experimental setups, demonstrating the possibility of running large models locally at the speed of interactive inference.
@@ Appendix A Equivalence of Optimal Tree Search to Shortest Path Search
Here, $x_{i}\in\tau$ refers to a token within the draft tree, $\mathcal{T}^{K}$ is a set of all trees of $K$ tokens and $P_{\text{reach}}(x_{i}|\tau)$ is the probability that the Speculative Execution verification phase accepts the full prefix $x_{0},\dots,x_{i-1}$ along the draft tree and considers sampling $x_{i}$ next. Finally, $P_{\text{accept}}(x_{i}|\tau)$ is the probability that the token $x_{i}$ will be accepted if it is reached during verification. Both $P_{\text{reach}}$ and $p_{\text{accept}}$ depend on the target model probabilities $P(x_{t+1}|x_{0:t},\theta_{\text{target}})$, which cannot be accessed in the drafting phase. Instead, we use the draft model to approximate the target model probabilities as follows:
@@
Since token probabilities cannot be greater than 1, the cumulative probability of $\pi(x_{i},\tau)\oplus x_{i}$ cannot exceed the cumulative probability of all tokens in $\pi(x_{i},\tau)$. Therefore, if a token $x_{i}$ is among the $K$ most likely tokens, every token in $\pi(x_{i},\tau)$ is also a part of the solution. Using this property, we can simplify Equation 3 as finding top-$K$ most likely prefixes, since they are guaranteed to form a tree. Formally speaking, the optimal tree consists of K tokens with the highest cumulative probability:
@@
This is similar (but not equivalent) to the standard beam search algorithm for neural sequence models Graves [2012], Boulanger-Lewandowski et al. [2013], Sutskever et al. [2014]. The main difference is that beam search looks for complete sequences, while we need a tree of partial drafts. However, using beam search instead of solving Equation 3 directly leads to suboptimal drafts (see Appendix F for details).
Instead, we solve Equation 4 by reformulating it as a special case of the shortest path search problem. More specifically,
@@
In the worst case, this algorithm still makes up to $K$ steps if the solution to Equation 3 is a single “stem” B tokens long. However, the actual number of steps is significantly lower, often slightly above the lower bound $\lceil B/K\rceil$. In the practical GPU implementation, we also limit the maximum depth with a parameter $D$. The purpose of D is to limit the edge case where the draft model is very confident about the next token, and thus the solution to Equation 3 is a single sequence of length K. For this edge case, Algorithm 2 will take long to generate a sequential draft, most of which will later be discarded if the draft model makes even one mistake.
Appendix B SpecExec Algorithm Diagram
Figure 6 displays a block diagram that outlines the key steps of the SpecExec algorithm.
[…]
Figure 6: A high-level overview of the SpecExec algorithm.
Appendix C Parallel Graph Search
@@
There are dozens of works that study practical implementations of parallel shortest path search and SSSP in particular. One line of work proposes inexact search algorithms that use an approximate priority queue to improve the performance of SSSP: Nguyen et al. [2013] proposes a queue with an integer metric, and Zhang et al. [2020] adds bucket fusion to reduce the synchronization overhead..
@@
A significant effort was dedicated to efficient shortest-path search on GPUs, among others. Harish and Narayanan [2007] proposes a GPU-efficient SSSP that outperforms sequential CPU computations. Davidson et al. [2014], Wang et al. [2016] compare several SSSP variants for GPUs. Iacono et al. [2019] adapts priority queues to run efficiently on GPU and uses the resulting data structure to accelerate SSSP.
Many works on parallel graph search focus on the distributed setting Malewicz et al. [2010], Zhu et al. [2016], Besta et al. [2017], addressing communication and synchronization overheads. Finally, a large body of work studies the theoretical properties of parallel SSSP, including Ullman and Yannakakis [1990], Klein and Subramanian [1997], Cohen [1993], Shi and Spencer [1999], Cohen [2000], Spencer [1997], Meyer [2001], Blelloch et al. [2016].
Appendix D Additional Implementation Details
Our system design follows the following loop:
  1. 1.
    Load the draft model onto GPU memory and generate a draft tree;
  2. 2.